OpenStudy (anonymous):

Mat obtains a mortgage for $250,000 with the following terms: - 15 year 5/1 ARM at 6.5% with a 3/7 cap structure - - Initial monthly payments: $2,177.77 What will be the balance of the loan at the end of the initial interest rate period? I got 191,792.80

OpenStudy (amistre64):

the key here is understanding the 5/1 notation. what do you recall?

OpenStudy (amistre64):

and if you have covered balloon payments, thats basically what its asking

OpenStudy (amistre64):

250000k^(5*12) - 2177.77(1-k^(5*12))/(1-k). k=(1+.065/12) http://www.wolframalpha.com/input/?i=250000k%5E%285*12%29+-+2177.77%281-k%5E%285*12%29%29%2F%281-k%29%2C+k%3D%281%2B.065%2F12%29 191 793 is fine

OpenStudy (anonymous):

Thank you. Can you check a couple more?

OpenStudy (amistre64):

if my sanity holds out sure :)

OpenStudy (anonymous):

Hahah cool, thank you.

OpenStudy (anonymous):

Marie has a 20 year adjustable rate mortgage with a fixed rate for the first 7 years. In the 8th year, the interest rate rises to 6.2%. the remaining balance at the end of the 7th years is $398,381.20. What is the monthly payment in the 8th year? I got 2553.73

OpenStudy (amistre64):

so we pretty much want to take out a loan for the remaining balance for the remaining time period at the adjusted rate .... correct?

OpenStudy (amistre64):

Bk^n(1-k)/(1-k^n) = P 398,381.20(1+.062/12)^(12(20-7))(1-(1+.062/12))/(1-(1+.062/12)^(12(20-7))) http://www.wolframalpha.com/input/?i=398%2C381.20%281%2B.062%2F12%29%5E%2812%2820-7%29%29%281-%281%2B.062%2F12%29%29%2F%281-%281%2B.062%2F12%29%5E%2812%2820-7%29%29%29 im getting something else, did you use the adjusted time frame? 20 - 7

OpenStudy (anonymous):

Hmm okay is it 177711.51?

OpenStudy (amistre64):

show me you work if you can

OpenStudy (amistre64):

you are most likely using a formula, what is your formula?

OpenStudy (amistre64):

and are there options?

OpenStudy (amistre64):

i use my own formula that i develped during the class simply becuase it is sooo useful in so many applications instead of trying to recall all the other formulas for each situation.

OpenStudy (amistre64):

a little workings .. we have a balance B, that gets interest ,r, added to it and then a payment ,P, removed: Bo = B B1 = Bo(1+r) - P B2 = B1(1+r) - P = Bo(1+r)^2 - P(1+r) - P B3 = B2(1+r) - P = Bo(1+r)^3 - P(1+r)^2 - P(1+r) - P .... let k clean up the compounding rate stuff to read easier Bn = Bok^n -P(1+k+k^2+...+k^(n-1)) since P represents a geometric series ... Bn = Bok^n - P(1-k^n)/(1-k) and thats the formula i use

OpenStudy (anonymous):

The other options are 119333.90 and 234027.44

OpenStudy (amistre64):

to determine monthly payment, then when Bn = 0, for n periods P = Bo k^n(1-k)/(1-k^n) n=12(20-7) k=1+.062/12 Bo = 398381.20

OpenStudy (anonymous):

Okay so 234027.44?

OpenStudy (amistre64):

im getting 3725.90

OpenStudy (anonymous):

Wait, I was looking at a different question! Oh man im sorry I gave the wrong options. but now it makes sense. Im so sorry

OpenStudy (amistre64):

youre options arent making any sense, ... :)

OpenStudy (anonymous):

Haha I know I see that too haha.

OpenStudy (anonymous):

Can you check a couple more?

OpenStudy (amistre64):

has this one been satisfied? we had different solutions, do we know why?

OpenStudy (anonymous):

Haha yes we do haha. Again im sorry for the different ones.

OpenStudy (amistre64):

lets do another then

OpenStudy (amistre64):

..... testing my sanity i spose lol

OpenStudy (anonymous):

Hahah thank you :)

OpenStudy (anonymous):

An investor obtains a balloon mortgage with the terms shown below: $370,00 20/5 balloon 5.1 % annual rate What is the monthly payment on this mortgage? I got $2,184.59

OpenStudy (amistre64):

monthly payment is the same as a 20 year fixed rate loan P = Bo k^n (1-k)/(1-k^n) P = 370000 k^(12*20) (1-k)/(1-k^(12*20)), k=(1+.051/12) P=2462.32 it would be helpful to know how you are approaching your solutions

OpenStudy (anonymous):

This one I guessed, but how did you get 2462.32?

OpenStudy (amistre64):

i posted how i got the formula i use and then just used the formula ...

OpenStudy (amistre64):

what are the options if any?

OpenStudy (anonymous):

The other ones are 6999.32, 6481.17, 2462.32, and the first one I did.

OpenStudy (amistre64):

id go with 2462 then :)

OpenStudy (amistre64):

theres a monthly payment formula in your course materials, my own formula is an unsimplified version of it that just makes more sense to me personally

OpenStudy (anonymous):

Oh okay cool then :) They should teach your formula. I have another question that is similar.

OpenStudy (amistre64):

lets try it out

OpenStudy (anonymous):

Balloon Mortgage: $300,000 25/5 balloon 6% annual rate Initial Monthly Payment $1932.90 What is the amount of the balloon payment rounded to the nearest dollar? I got $271,145

OpenStudy (anonymous):

Or is it $280,559?

OpenStudy (amistre64):

so the remaining balance after 5 years of payments, 5*12 = 60 payment periods B60 = Bo k^(60) -P(1-k^(60))/(1-k) B60 = 300000 k^(60) -1932.90(1-k^(60))/(1-k), k=1+.06/12 http://www.wolframalpha.com/input/?i=300000+k%5E%2860%29+-1932.90%281-k%5E%2860%29%29%2F%281-k%29%2C+k%3D1%2B.06%2F12 about 269 797 is what i get

OpenStudy (anonymous):

That's one of the choices. How come its not 271145?